OpenStudy (anonymous):

Is this correct? 45-5x=5x+55 doesn't x = 1O ?

OpenStudy (skullpatrol):

Nope,

OpenStudy (anonymous):

wouldn't x =-1

OpenStudy (anonymous):



OpenStudy (anonymous):

 move 5x over, and then move 45 over

OpenStudy (anonymous):

what i did : 45-5x=5x+55 55-45=10 -5x-5x= x x=10

OpenStudy (anonymous):

you would get -10x=10

OpenStudy (anonymous):

thats right , thank you :)

OpenStudy (anonymous):

after that you would divide by -10 on both sides and x would equal -1

OpenStudy (anonymous):

your welcome :)

OpenStudy (skullpatrol):

@_Meriah_ you can always check your own answer by substituting that value back into the original equation.

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

yes i have done that with some of my equations , thank you :)

OpenStudy (skullpatrol):

Thanks for trying to understand :)